# A Survey of Cross-Lingual Alignment: Definitions, Methods, Future Understanding Cross-Lingual Alignment—A Survey

### v1

@@ B.2 What to use?
Results on zero-shot transfer overall show a similar picture to XNLI, although details change. For example, XLM-ALIGN’s performance stands out on UD-POS but is “only” competitive on the other tasks. HiCTL, meanwhile, is fairly competitive in zero-shot XNLI performance but falls a bit further behind in Table 2. The authors of mDeBERTaV3 do not report any of these other tasks, leaving XLM-ALIGN, XLM-E, and ERNIE-M to take the top spots: they all perform well on these three tasks but alternately take the lead.
In the translate-train setting (Table 3), VECO performs best on all three tasks, with HiCTL on par for PAWS-X but not UD-POS or MLQA. For XNLI, the best translate-train performance was attained by X-MIXUP, which still does well on these tasks. Again, overall trends are very similar as for the XNLI task.
@@ B.3 Limitations
| InfoXLM | 270M | 76.5 | – | – | 68.1 |
| ERNIE-M | 270M | 77.3 | – | – | 68.7 |
| HiCTL | 270M | 77.3 | 71.4 | 84.5 | 65.8 |
@@
| HiCTL | 550M | 81.0 | 74.8 | 87.5 | 72.8 |
| InfoXLM | 550M | 81.4 | – | – | 73.6 |
| ERNIE-M | 550M | 82.0 | – | 89.5 | 73.7 |
| XLM-R + xTune | 550M | 82.6 | 78.5 | 89.8 | 74.4 |
@@
| ERNIE-M | 550M | 84.2 | – | 91.8 | – |
| HiCTL | 550M | 84.5 | 76.8 | 92.8 | 74.4 |
| XLM-R + X-MIXUP | 550M | 85.3 | 78.4 | 91.8 | 76.5 |
| mT5 | 580M | 79.8 | – | 89.3 | 75.3 |
| VECO | 662M | 84.3 | 79.8 | 92.8 | 77.5 |
@@
| ERNIE-M | 550M | 87.9 | – |
| LAPCA-LM | 550M | – | 83.5 |

### v2

@@ B.2 What to use?
Results on zero-shot transfer overall show a similar picture to XNLI, although details change. For example, XLM-ALIGN’s performance stands out on UD-POS but is “only” competitive on the other tasks. HiCTL, meanwhile, is fairly competitive in zero-shot XNLI performance but falls a bit further behind in Table 2. The authors of mDeBERTaV3 do not report any of these other tasks, leaving XLM-ALIGN, XLM-Ebase, and ERNIE-Mbase to take the top spots: they all perform well on these three tasks but alternately take the lead.
In the translate-train setting (Table 3), VECOin performs best on all three tasks, with HiCTLlarge on par for PAWS-X but not UD-POS or MLQA. For XNLI, the best translate-train performance was attained by X-MIXUP, which still does well on these tasks. Again, overall trends are very similar as for the XNLI task.
@@ B.3 Limitations
InfoXLMbase | 270M | 76.5 | – | – | 68.1 |
ERNIE-Mbase | 270M | 77.3 | – | – | 68.7 |
HiCTLbase | 270M | 77.3 | 71.4 | 84.5 | 65.8 |
@@
HiCTLlarge | 550M | 81.0 | 74.8 | 87.5 | 72.8 |
InfoXLMlarge | 550M | 81.4 | – | – | 73.6 |
ERNIE-Mlarge | 550M | 82.0 | – | 89.5 | 73.7 |
XLM-Rlarge + xTune | 550M | 82.6 | 78.5 | 89.8 | 74.4 |
@@
ERNIE-Mlarge | 550M | 84.2 | – | 91.8 | – |
HiCTLlarge | 550M | 84.5 | 76.8 | 92.8 | 74.4 |
XLM-Rlarge + X-MIXUP | 550M | 85.3 | 78.4 | 91.8 | 76.5 |
mT5base | 580M | 79.8 | – | 89.3 | 75.3 |
VECOin | 662M | 84.3 | 79.8 | 92.8 | 77.5 |
@@
ERNIE-Mlarge | 550M | 87.9 | – |
LAPCA-LMlarge | 550M | – | 83.5 |